OpenStudy (anonymous):

Willl give medal!!! Please help! 1.Write the equation of the line which passes through (5, –2) and is parallel to x = 4. (1 point) 2.Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2. (1 point) 3.Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. (2 point) 4.Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form. (2 point)

OpenStudy (anonymous):

1. x=5 2, y=1 3. (y-2)=-(x+4) y=-x-2 4. (y+3)=-1/4 (x-2) y=-1/4 x + 1/2 - 3 y=-1/4 x - 5/2

OpenStudy (anonymous):

you can check it this graphing tool which is pretty cool https://www.desmos.com/calculator